OpenStudy (anonymous):

Find the values of c and d that make the following function f(x)=\left\lbrace \begin{array}{cl} 2 x &\mbox{if } x<1\\ cx^2+d &\mbox{if } 1\le x <2\\ 5 x &\mbox{if } x\ge 2\\ \end{array}\right.

OpenStudy (anonymous):

A moment of reflection reveals that 2(1)=1^2 c+d 2^2 c +d = 5(2) so c+d=3 4 c + d =10 Solve for c and d

OpenStudy (anonymous):

c+d =2 (Sorry)

OpenStudy (anonymous):

\[ \left\{\left\{c\to \frac{8}{3},d\to -\frac{2}{3}\right\}\right\} \]

OpenStudy (anonymous):

^hmm interesting

OpenStudy (anonymous):

I wonder how you come to this conclusion 2(1)=1^2c+d

OpenStudy (anonymous):

because it was clearly stated that f(x)=2x when x<1 and need not to be equal to 1 please correct if I'm wrong

OpenStudy (anonymous):

The limit from both sides should be equal

OpenStudy (anonymous):

I cannot see any evidents saying the function is continuos